Subject: Minutes of March 11, 1996 APO User's Committee Meeting

From: Michael Strauss

Submitted: Mon, 11 Mar 1996 20:26:58 -0500 (EST)

Message number: 53 (previous: 52, next: 54 up: Index)

  Minutes of APO 3.5m User's Committee Phone Conference
		March 11, 1996

  Attending: Alan Uomoto, Bruce Gillespie, Ed Turner, Lew Hobbs,
Michael Strauss, Rene Walterbos, Alan Uomoto, Julie Lutz, Alan Diercks

  Dormitory Policy.
  Observing Specialists and their role in monitoring weather
  Status update since last month.
  Enclosure Wheel. 
  Throughput issues.
  Instrument engineer.
  Discussion of secondary.
  How to structure these meetings.

Dorm Policy: People should file reservations at least 5 days in
advance. This is necessary in order to schedule the housekeeper
properly. Late reservations may be refused (you will be supplied with
a list of local hotels where you may make your own arrangements) if no
cleaned rooms are available and if there is no legitimate reasons the
reservations could not have been filed on time.

Observing specialists have the responsibility to decide when to close the
dome. Please do not challenge or argue their decisions; also, once
the dome has been closed, they will closely monitor conditions and keep
you informed; don't bother them by pinging them every few minutes.  See
the criteria for dome closure on the observatory web site.  

Institutions are charged when the telescope time is lost due to bad
weather. For a future meeting, should we modify this charging system?

Status report:
  Finishing 1 week shutdown. Tightened up vanes holding up secondary
to try to get rid of some harmonic motions (10 Hz and 20 Hz; sound
familiar?). So this is believed to be one of the primary sources of
vibration (WIYN may have similar problems). The telescope has been
recollimated. Some hints that the focal plane may be tilted relative
to the instrument rotation plane. May require adjustment of the
tertiary, to fix this. 
  Bruce believes that half of the cause of the poorer seeing is due to
various vibrations, and half due to the nature of the secondary. The
process of pinning down these vibrations is continuing; shall we
pre-empt more engineering time for this?

  The worst cracked wheel is to be replaced over the next few days. It
should be done by Wednesday, but it might take longer! The other
wheels (in much better shape) will be replaced over the next year or
so. It may be that the original wheels were never flame-hardened.

  The DIS apparently has a narrower slit than stated; it is 1.0
arcsec, not 1.5 arcsec wide, according to Karen Gloria (apo35-dis
message #11). 

Throughput of telescope:
  There have been two recent, independent determinations of the
throughput of the DSC, by Krys Stanek and Xiao-Hui Fan. Numbers are
~15% in g and r, very similar to numbers from May '95 by Richmond and
Stanek. Why is it so low?
  There is also anecdotal evidence that the DIS has low throughput,
especially in the red. Michael Strauss reports taking spectra of red
giants, and having the blue saturate *well* before the red. The total
throughput is such that "it doesn't feel like observing on the 3.5m
telescope". Jon Davis and Jon Brinkmann will have the job of measuring
and characterizing telescope and instrument behavior.

Now have two instrument engineers. Jon Brinkman is spending a few
months in Princeton working on the SDSS camera with Jim Gunn. There is
An engineer named Steve Knapp will arrive at the site in a week or two
for three to six months or so to take Jon's place while he is
  The question of the overall throughput of the telescope and
instruments is high priority!

The secondary and the image quality:
  If the seeing is getting as good as 0.7", does that weaken the case
for getting a new secondary? Is the secondary much better than we
think? It remains somewhat ambiguous whether the secondary is indeed
the cause of the poor seeing. The ultimate test would be a prime-focus
Hartman test. Bruce argues for a new secondary truss to carry out this
test (of the order of $12K). Another possibility would be to do an
full-aperture test at one of several labs. This would be much more
detailed than would be the profilometry that already exists. This
would have the advantage of being quick; it would take all of a
week. It would cost ~$5K.
  Bruce will look into these various options. 
  Question: is the poor seeing all simply a problem of poor
collimation? Answer: Apparently not.

  Schedule has two week-long engineering shutdowns; one of these
could be used to do this test.

  One other test would be to rotate secondary by 120 degrees. This
would be of the order of a week. 

  Schedule is about to come out. Engineering time next quarter is 20%
(was 28% last quarter, 40% quarter before that).

  How best to run these meetings? Should we have a written progress report
once per meeting? It puts time pressure on Bruce. Perhaps Bruce
should put together a list of concerns once a month; in a sense,
suggest agenda items for discussion. Rene: the members of the user's
committee should bring up those issues that concern the users.

  All committee members should test their SLIP connections at their
home institutions. 

  Next meeting, Monday, April 8, 12 PM EST. 
APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 53 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to